Stock Investment Transactions

On September 12, 2,900 shares of Aspen Company are acquired at a price of $36.00 per share plus a $145 brokerage commission. On October 15, a $1.20-per-share dividend was received on the Aspen Company stock. On November 10, 1,160.00 shares of the Aspen Company stock were sold for $31 per share less a $58 brokerage commission.

When required, round final answers to the nearest dollar. For a compound transaction, if an amount box does not require an entry, leave it blank.

Prepare the journal entries for the original purchase, the dividend, and the sale under the cost method.

Debit . Credit

Sept. 12
Oct. 15
Nov. 10
Answer #1
Sept 12 Investments-Aspen Company stock 104545 =(2900*36)+145
              Cash 104545
Oct. 15 Cash 3480 =2900*1.20
          Dividend Revenue 3480
Nov. 10 Cash 35902 =(1160*31)-58
Loss on sale of investments 5916
         Investments-Aspen Company stock 41818 =104545*1160/2900